# Pareto-optimality of pulses for robust population transfer in a ladder-type qutrit

### v2

@@ Abstract
Frequency-modulation schemes have recently emerged as alternatives to standard Rabi pulses for realizing robust quantum operations. In this work, we investigate short-duration population transfer between the ground and first excited states of a ladder-type qutrit, with the goal of minimizing leakage into the second excited state. Our multiobjective approach seeks to reduce the maximum transient second-state population and maximize detuning robustness. Inspired by two-state models—such as the Allen-Eberly and Hioe-Carroll models—we extend these concepts to our system, exploring a range of pulse families, including those with super-Gaussian envelopes and polynomial detuning functions. We identify Pareto fronts for pulse models constructed from one of two envelope functions paired with one of four detuning functions. We then analyze how each Pareto-optimal pulse parameter influences the two Pareto objectives and amplitude robustness.
pacs:
Valid PACS appear here
I Introduction
@@
Robust population transfer remains a key area of interest in quantum control. Various techniques employing adiabatic passage have been the primary means to this end, engineered to deliver high fidelity and efficiency with minimal dissipation. The principle behind these methods lies in the concept of adiabaticity, which ensures that a system remains in its instantaneous eigenstate throughout its evolution. For example, in a two-level system with Hamiltonian $\hat{H}=\frac{\hbar}{2}(\Delta\hat{\sigma}_{x}+\Omega\hat{\sigma}_{z})$, the energy eigenvalues are $E_{\pm}=\pm\frac{\hbar}{2}\sqrt{\Delta^{2}+\Omega^{2}}$, corresponding to eigenstates $|-\rangle=\frac{1}{\sqrt{2}}(\cos(\theta/2)|0\rangle-\sin(\theta/2)|1\rangle)$ and $|+\rangle=\frac{1}{\sqrt{2}}(\sin(\theta/2)|0\rangle+\cos(\theta/2)|1\rangle)$, where $\theta$ is the so-called mixing angle and relates to the amplitude and detuning via $\tan(\theta)=\Omega/\Delta$. The system evolves adiabatically if $\dot{\theta}/\Delta E\ll 1$, which is satisfied when $\frac{\norm{\dot{\Omega}\Delta-\Omega\dot{\Delta}}}{(\Omega^{2}+\Delta^{2})^{3/2}}\ll 1$ [1]. When this condition is met, population transfer occurs through gradual evolution along eigenstates, avoiding unwanted transitions.
@@
Rapid adiabatic passage (RAP) is the staple technique for robust population inversion. RAP was first realized in NMR, where the spins on the magnetic moments are inverted by sweeping the drive field frequency across the resonance [2]. It is rapid in the sense that the frequency sweep is faster than the relaxation time of the spins, or in the case of circuit QED, the relaxation time of the qubit. RAP is very versatile in that it can be implemented on any two-level quantum system with an external drive field. As such, it has found many applications in atomic, molecular, and optical physics [3, 4, 5, 6], as well as in chemical reactions [7], interferometry [8, 9], and circuit QED [10, 11]. In circuit QED, the cited works utilize Stark-chirped rapid adiabatic passage (SCRAP), a variety of RAP in which a slowly varying Stark shift sweeps the energy of the target state through resonance to induce complete population transfer [12].
In the context of RAP, pulse shaping is essential for ensuring efficient and robust population transfer. Gaussian pulses are often employed [13, 14], in part because they have been shown to reduce leakage beyond the computational space [15]. Super-Gaussian pulses have also been utilized to enhance control over transition dynamics. For instance, our previous work on high-fidelity qubit control using phase-modulated pulses demonstrated that super-Gaussian envelopes can suppress amplitude more effectively compared to their Gaussian counterparts, providing improved robustness against parameter variations [16]. Since the phase modulation we employ is equivalent to a frequency modulation through the choice of a frame rotation, we hereafter refer the scheme as one that utilizes frequency modulation.
@@
A central goal in quantum control of weakly anharmonic systems, such as the transmon, is to suppress leakage outside the computational subspace—that is, to confine population dynamics to the lowest two energy levels. When driving the ground-to-first-excited-state transition $\mathrm{g-e}$ with a resonant microwave pulse, the close proximity of $\mathrm{e-f}$ transition frequency means that higher levels cannot be ignored. The spectral width of the pulse—particularly if it is short and strong—can unintentionally drive the $\mathrm{e-f}$ transition, leading to population leakage out of the computational subspace. To mitigate this, techniques such as derivative removal by adiabatic gate (DRAG) pulses are used, which shape the drive to cancel unwanted transitions [17]. Alternatively, using longer pulses reduces spectral overlap but can increase sensitivity to decoherence [18]. Ultimately, achieving high-fidelity control in such systems requires carefully balancing gate speed, spectral selectivity, and pulse shaping to minimize leakage while maintaining fast, accurate operations [19, 20].
@@
In this work, we pursue this goal by building on rapid adiabatic passage (RAP) methods and our aforementioned work, where the drive frequency is modulated, in the simplest case, by adding to the qubit frequency an instantaneous linear detuning $\Delta(t)=2\Delta_{\rm{max}}t/T$, parametrized in terms of modulation depth $\Delta_{\rm{max}}$ and pulse duration $T$, with $t\in[-T/2,T/2]$. Furthermore, following our frequency-modulation technique, we employ a single continuously modulated pulse and explore different RAP-like models on a three-level ladder system with states $\{|\rm{g}\rangle,|\rm{e}\rangle,|\rm{f}\rangle\}$, and seek to minimize leakage into the second excited state $|\rm{f}\rangle$ when driving population from the ground state to the first excited state. Specifically, we wish to minimize the maximum transient second excited state population $\max_{t}(p_{\rm{f}})$, with $t\in[-T/2,T/2]$, and simultaneously maximize the detuning robustness at a threshold of the final first excited state population $p_{\rm{e}}(T/2)\geq 0.99$. As this is a multiobjective optimization problem, we seek to find the so-called Pareto front which meets our criteria. A Pareto front is a solution to the multiobjective optimization in which it proves impossible to improve one objective without making the other worse. Pareto fronts have found applications in a plethora of disciplines including economics [21], traffic regulation [22], machine learning [23], and bioinformatics [24]. In physics, Pareto’s methodology has been employed to find the extrema of observables as a form of optimal quantum control [25, 26].
@@
This paper is organized as follows: In Section II we describe the methodology for our frequency-modulated robust population transfer scheme, introduce the pulse models used in this work, and detail the underlying methodology for the Pareto optimization. In Section III, we present the Pareto fronts for the different models, each at two pulse durations, and assign an effective limit to the fronts by utilizing the Landau-Zener formula. We perform parameter analysis in Section IV, and ascertain the most impactful parameters for each of our objective outcomes. Finally, we summarize the key findings of this work in Section V.
II Methodology
We employ a single, continuously modulated drive pulse with Rabi amplitude $\Omega(t)$ and detuning $\Delta(t)$ corresponding to the first transition in order to emulate rapid adiabatic passage between the ground state $|\rm{g}\rangle$ and the first excited state $|\rm{e}\rangle$. For brevity, we hereafter drop the subscript for the Rabi amplitude and detuning function, except when presenting the system Hamiltonian in the next subsection. The shape of the drive pulse plays a central role in achieving efficient and robust state transfer. As we describe in the subsection after next, we focus on two flat-at-the-top envelope functions that are known to deliver strong detuning robustness.
II.1 The frequency-modulation scheme adopted to a ladder qutrit
@@
An energy diagram of the modulation scheme is shown in Fig. 1. The diagonal part of the Hamiltonian in Eq. 1 corresponds to a ladder-type qutrit with energy level separations $\omega_{\rm{ge}}$ and $\omega_{\rm{ef}}$, in the frame co-rotating with the drive frequency. For a transmon qutrit, the ratio of the Josephson energy $E_{J}$ to the charging energy $E_{C}$ is very large, and in the extreme limit $E_{J}/E_{C}\rightarrow\infty$, the transmon energy levels approach those of a harmonic oscillator for low-energy states, with nearly constant but slightly decreasing spacing due to residual anharmonicity. The qutrit is driven by a microwave field with frequency $\omega_{d}(t)$.
@@
We consider a rotating frame such that for the ground state $|\rm{g}\rangle$ the energy is shifted by $-\Delta(t)$, and the zero of the energy is set at the first excited state $|\rm{e}\rangle$, where $\Delta(t)=\omega_{\rm ge}-\omega_{d}-(d\omega_{d}/dt)t$. The charging energy $E_{C}$ gives rise to the anharmonicity in the energy levels in superconducting qutrits such as the transmon, and relates to the transition frequencies via $E_{\rm C}=\hbar\omega_{\rm{ge}}-\hbar\omega_{\rm{ef}}$. Thus, the detuning of the $|\mathrm{f}\rangle$ level is $\Delta(t)-E_{\rm C}/\hbar$. The ratio of the Rabi frequencies of the two transitions is given by the ratio of the transition electrical dipole moments. The dipole operator is proportional to the number operator $\hat{n}=-i\partial/\partial\varphi$, and thus the dipole matrix elements are $g_{n,n+m}\sim\langle n|\hat{n}|n+m\rangle$, and consequently, $g_{12}/g_{01}=\sqrt{2}$ and $g_{02}=0$.
@@
Combining these considerations, we arrive at the Hamiltonian in the rotating wave approximation, as detailed in Ref. [27]:
@@
Figure 1: The energy diagram of our frequency modulation scheme shown for two different detuning functions, one linear (blue), and one quintic (green), for time $t\in[-T/2,T/2]$. The transition frequencies are $\omega_{\rm{ge}}$ and $\omega_{\rm{ef}}$, the modulation depth is denoted by $\Delta_{\rm max}$, and the frequency offset is $\delta$. The two-photon transition frequency $\omega_{\rm{gf}}^{\rm 2ph}$ is exactly halfway between $\omega_{\rm{ge}}$ and $\omega_{\rm{ef}}$.
In the following subsection, we discuss the Pareto-optimization technique for finding parameters that, when inserted into the matrix elements of Eq. 1, guide the system to adiabatically follow its lowest eigenstate, thereby facilitating robust population transfer from $|\rm{g}\rangle$ to $|\rm{e}\rangle$.
II.2 Pulse models
@@
Several envelope functions have been explored in two-state and three-state quantum models, each with distinct advantages and limitations [28, 29, 30, 31, 32, 33, 34, 14, 35]. Constant-amplitude pulses can drive transitions via Rabi oscillations but require precise timing and are highly sensitive to detuning errors [36]. Landau-Zener transitions, which use a linearly time-dependent detuning with a constant coupling, offer analytical solutions but can suffer from unwanted population leakage [37, 38]. Further, exponential pulses have been applied in time-dependent quantum transition models but often suffer from long tails, reducing efficiency by causing non-negligible residual population transfer outside the intended time window [39, 40].
@@
Smoothly varying pulses, such as Gaussian and hyperbolic secant envelopes, provide greater control over transition probabilities. Gaussian pulses minimize spectral side lobes and have been widely used in adiabatic passage schemes [13, 41, 42, 43, 14, 35]. However, achieving complete population transfer with Gaussian pulses requires precise detuning control. On the other hand, hyperbolic secant pulses are analytically solvable and naturally suited for robust adiabatic and nonadiabatic population transfer, making them widely used in RAP and related techniques [28, 29, 30, 31, 32, 44, 45].
@@
In this work, we focus on two pulse envelopes: generalized super-Gaussian and hyperbolic secant varieties. Both envelope types are flat at the top, meaning the pulse amplitude remains approximately constant near the center of the drive, around $t=0$, where the avoided crossing typically occurs. This flat-top structure grants a greater degree of detuning robustness by maintaining a high Rabi frequency over the critical transition region, thereby extending the adiabatic window and improving population transfer even under moderate detuning errors [16]. The super-Gaussian envelope allows for greater control over the transition window by flattening the peak and steepening the decay, which can simultaneously be advantageous in terms of requiring lowering amplitudes. Meanwhile, the hyperbolic secant envelope serves as a benchmark due to its well-established analytical properties. These envelope functions are defined as follows:
@@
for the super-Gaussian envelope, and
@@
By employing these two pulse shapes, we aim to explore their effects on state transfer efficiency and robustness, particularly in the context of our frequency-modulation scheme adopted to a three-level system. Each of these envelopes contain three parameters which we seek to multiobjectively optimize with respect to the minimization of the maximum second excited state population $\max_{t}(p_{\rm{f}})$ reached across the pulse duration $t\in[-T/2,T/2]$, and (simultaneously) with respect to the detuning robustness defined as the frequency width of the first excited state population at the threshold $p_{\rm{e}}(T/2)\geq 0.99$ with a local maximum of $p_{\rm{e}}(T/2)\geq 0.999$ within this width. The envelope parameters are amplitude $\Omega_{0}$, envelope order $n$, and the parameter $\beta$ (for super-Gaussian pulses) or $\epsilon$ (for hyperbolic secant pulses), which determine the envelope width and steepness. The variable order $n$ generalizes the super-Gaussian envelopes in that if $n=2$ we recover a Gaussian profile, $n=4$ corresponds to the so-called super-Gaussian variety, and as $n\rightarrow\infty$ the pulse approaches a rectangular form. Amplitudes $\Omega_{0}$ are searched for in the range $[0,0.1]$ GHz, consistent with experimental constraints that help mitigate parasitic effects such as leakage into $|\rm{f}\rangle$ and higher states, as well as AC Stark shifts, which are known to degrade qubit coherence and fidelity in similar systems [46]. The envelope order $n$ is scanned over the range $(0,10]$, while $\beta$ and $\epsilon$ are searched within $[-6,0)$ and $(0,6]$, respectively.
Additionally, we consider four detuning models and assign each to the two envelope types. Our first detuning model is linear, and the model name when assigned to the super-Gaussian envelope is SG1, and SECH1 when assigned to the hyperbolic secant envelope. For subsequent models, we hereafter adopt the prefixes SG and SECH for the model names. The modulation depth $\Delta_{\rm max}$ has an upper limit of half the charging energy, which is set at $E_{C}/2=0.15$ GHz in this work, since it is the maximum depth over which we can sweep in a typical circuit QED experiment using a transmon.
@@
where the prefactor $k_{1}$ determines the slope of our detuning function and is a parameter we optimize by searching in the range $k_{1}\in[-0.3,0.3]$ GHz. Since Eq. 4 is a linear detuning function, the parameter $k_{1}=\Delta_{\rm{max}}$. Further, we center our modulation about the frequency offset $\delta$, i.e., by sweeping $(2\Delta_{\rm max}/T)t$ from $t=-T/2$ to $t=T/2$.
We also explore a quintic detuning function, i.e.,
$\Delta(t)=\sum\limits_{n=1}^{5}k_{n}(2t/T)^{n}+\delta,$ (5)
where $k_{n>1}$ are each searched for in the range $[-0.1,0.1]$ GHz. For this detuning function, the effect of the parameter $k_{1}$ is the same as in Eq. 4; whereas, the prefactor for the quadratic term, $k_{2}$, makes the detuning function convex (upward-bending) when $k_{2}>0$ and concave (downward-bending) when $k_{2}<0$. Further, the parameter $k_{3}$, the cubic prefactor, introduces asymmetry and has the effect of making the right side ($t>0$) of the detuning function steeper and the left side ($t<0$) shallower when $k_{3}>0$, and vice versa when $k_{3}<0$. Still further, the quartic ($k_{4}$) and quintic ($k_{5}$) prefactors have similar effects to $k_{2}$ and $k_{3}$, respectively, but with stronger effects at larger |$\absolutevalue{t}$ |. Overall, the symmetry properties are governed by the parity of each term: the even-powered terms ($k_{2}$, $k_{4}$) contribute symmetric components about $t=0$, while the odd-powered terms ($k_{1}$, $k_{3}$, $k_{5}$) contribute antisymmetric components. A detuning function composed purely of odd-powered terms is antisymmetric about the origin, but adding even-powered terms breaks this antisymmetry, resulting in an overall asymmetric function. Thus, the inclusion of both even and odd terms enables finer control over both the shape and symmetry of the detuning profile.
We also examine detuning functions that incorporate the hyperbolic tangent. Motivated by the Allen–Eberly [31] and Hioe–Carroll [32] models—both of which employ hyperbolic secant pulse envelopes—we respectively consider the following forms:
$\Delta(t)=k_{1}\tanh(2\gamma_{1}t/T)+\delta,$ (6)

### v4

@@ Abstract
Frequency-modulation schemes offer an alternative to standard Rabi pulses for realizing robust quantum operations. In this work, we investigate short-duration population transfer between the ground and first excited states of a ladder-type qutrit, with the goal of minimizing leakage into the second excited state. Our multiobjective approach seeks to reduce the maximum transient second-state population and maximize detuning robustness. Inspired by two-state models—such as the Allen-Eberly and Hioe-Carroll models—we extend these concepts to our system, exploring a range of pulse families, including those with super-Gaussian envelopes and polynomial detuning functions. We identify Pareto fronts for pulse models constructed from one of two envelope functions paired with one of four detuning functions. We then analyze how each Pareto-optimal pulse parameter influences the two Pareto objectives as well as amplitude robustness.
keywords:
Quantum control, Pareto front, Robust population transfer, Rapid adiabatic passage, Frequency modulated pulses
1 Introduction
@@
Robust population transfer remains a key area of interest in quantum control. Various techniques employing adiabatic passage have been the primary means to this end, engineered to deliver high fidelity and efficiency with minimal dissipation. The principle behind these methods lies in the concept of adiabaticity, which ensures that a system remains in its instantaneous eigenstate throughout its evolution. For example, in a two-level system with Hamiltonian $\hat{H}=\frac{\hbar}{2}(\Omega\hat{\sigma}_{x}+\Delta\hat{\sigma}_{z})$, the energy eigenvalues are $E_{\pm}=\pm\frac{\hbar}{2}\sqrt{\Delta^{2}+\Omega^{2}}$, corresponding to eigenstates $|-\rangle=\cos(\theta/2)|0\rangle-\sin(\theta/2)|1\rangle$ and $|+\rangle=\sin(\theta/2)|0\rangle+\cos(\theta/2)|1\rangle$, where $\theta$ is the so-called mixing angle and relates to the amplitude and detuning via $\tan(\theta)=\Omega/\Delta$. The system evolves adiabatically if $\dot{\theta}/\Delta E\ll 1$, which is satisfied when $\frac{\norm{\dot{\Omega}\Delta-\Omega\dot{\Delta}}}{(\Omega^{2}+\Delta^{2})^{3/2}}\ll 1$ [Vitanov_2001]. When this condition is met, population transfer occurs through gradual evolution along eigenstates, avoiding unwanted transitions.
@@
Rapid adiabatic passage (RAP) is the staple technique for robust population inversion. RAP was first realized in NMR, where the spins on the magnetic moments are inverted by sweeping the drive field frequency across the resonance [Bloch1946]. It is rapid in the sense that the frequency sweep is faster than the relaxation time of the spins, or in the case of circuit QED, the relaxation time of the qubit. RAP is very versatile in that it can be implemented on any two-level quantum system with an external drive field. As such, it has found many applications in atomic, molecular, and optical physics [Loy1974, Grischkowsky1976, Caussyn1994, Noel2012], as well as in chemical reactions [Kral2007], interferometry [Weitz1994, Kovachy2012], and circuit QED [Wei2008, Nie2010]. In circuit QED, the cited works utilize Stark-chirped rapid adiabatic passage (SCRAP), a variety of RAP, in which a slowly varying Stark shift sweeps the energy of the target state through resonance to induce complete population transfer [Yatsenko1999].
Beyond these implementations, RAP has also inspired a number of fast and robust quantum control protocols. For example, the Roland–Cerf protocol introduced a locally adiabatic schedule that slows near the avoided crossing to maintain high fidelity while reducing overall evolution time [Roland_2002]. This idea has since been experimentally realized and extended in various contexts, including high-fidelity quantum driving [Bason_2012], as well as resonant shortcuts to adiabatic passage using only longitudinal ($z$-field) control [Stefanatos_2019]. Similarly, RAP-inspired strategies have been employed to implement qubit gates using longitudinal ($\sigma_{z}$) control alone [Martinis_2014].
In the context of RAP, pulse shaping is essential for ensuring efficient and robust population transfer. Gaussian pulses are often employed [Demirplak_2003, Rangelov_2012], in part because they have been shown to reduce leakage beyond the computational space [Steffen_2003]. Super-Gaussian pulses have also been utilized to enhance control over transition dynamics. For instance, their flatter tops not only suppress spectral leakage and limit off-resonant excitations but also enhance detuning robustness, particularly near the pulse center—an effect especially beneficial for adiabatic or RAP-like protocols [Kuzmanovic_2024].
@@
A central goal in quantum control of weakly anharmonic systems, such as the transmon, is to suppress leakage outside the computational subspace—that is, to confine the population dynamics to the lowest two energy levels. When driving the ground-to-first-excited-state transition $\mathrm{g-e}$ with a resonant microwave pulse, the close proximity of the $\mathrm{e-f}$ transition frequency means that higher levels cannot be ignored. The spectral width of the pulse—particularly if it is short and strong—can unintentionally drive the $\mathrm{e-f}$ transition, leading to population leakage out of the computational subspace. To mitigate this, techniques such as derivative removal by adiabatic gate (DRAG) pulses are used, which shape the drive to cancel unwanted transitions [Motzoi_2009, Hyyppa_2024]. Alternatively, using longer pulses reduces spectral overlap but can increase sensitivity to decoherence [Khodjasteh_2011]. Ultimately, achieving high-fidelity control in such systems requires carefully balancing gate speed, spectral selectivity, and pulse shaping to minimize leakage while maintaining fast, accurate operations [Werninghaus_2021, Babu_2021].
@@
In this work, we pursue this goal by building on rapid adiabatic passage (RAP) methods and our previous work involving high-fidelity qubit control using phase-modulated pulses [Kuzmanovic_2024]. Since the phase modulation we employ is equivalent to a frequency modulation through an appropriate choice of frame rotation, we hereafter refer to the scheme as one that utilizes frequency modulation. In the simplest case, this drive frequency is modulated by adding to the qubit frequency an instantaneous linear detuning $\Delta(t)=2\Delta_{\rm{max}}t/T$, parametrized in terms of modulation depth $\Delta_{\rm{max}}$ and pulse duration $T$, with $t\in[-T/2,T/2]$. Furthermore, following our frequency-modulation technique, we employ a single continuously modulated pulse and explore different RAP-like models on a three-level ladder system with states $\{|\rm{g}\rangle,|\rm{e}\rangle,|\rm{f}\rangle\}$, and seek to minimize leakage into the second excited state $|\rm{f}\rangle$ when driving population from the ground state to the first excited state. Specifically, we wish to minimize the maximum transient second excited state population $\max_{t}(p_{\rm{f}})$, with $t\in[-T/2,T/2]$, and simultaneously maximize the detuning robustness at a threshold of the final first excited state population $p_{\rm{e}}(T/2)\geq 0.99$. As this is a multiobjective optimization problem, we seek to find the so-called Pareto front which meets our criteria. A Pareto front is a solution to the multiobjective optimization in which it proves impossible to improve one objective without making the other worse. Pareto fronts have found applications in a plethora of disciplines including economics [Newberry1984], traffic regulation [Jiao2016], machine learning [Jin2008], and bioinformatics [Gonzalez-Sanchez2023]. In physics, Pareto’s methodology has been employed to find the extrema of observables as a form of optimal quantum control [Chakrabarti2008, Sun2017].
@@
This paper is organized as follows: In Sect. 2 we describe the methodology for our frequency-modulated robust population transfer scheme, introduce the pulse models used in this work, and detail the underlying methodology for the Pareto optimization. In Sect. 3, we present the Pareto fronts for the different models, each at two pulse durations, and assign an effective limit to the fronts by utilizing the Landau-Zener formula. We perform parameter analysis in Sect. 4, and ascertain the most impactful parameters for each of our objective outcomes. Finally, we summarize the key findings of this work in Sect. 5.
2 Methodology
This section has three objectives. First, we describe the frequency modulation scheme adopted in this work. Second, we present the pulse models explored, along with the parameters targeted for optimization. Finally, we describe the algorithm employed to compute the Pareto front.
2.1 The frequency-modulation scheme adopted to a ladder qutrit
@@
We employ a single, continuously modulated drive pulse with Rabi amplitude $\Omega(t)$ and detuning $\Delta(t)$ with respect to the first transition in order to emulate rapid adiabatic passage between the ground state $|\rm{g}\rangle$ and the first excited state $|\rm{e}\rangle$. An energy diagram of the modulation scheme is shown in Fig. 1. The diagonal part of the Hamiltonian in Eq. (1) corresponds to a ladder-type qutrit with energy level separations $\omega_{\rm{ge}}$ and $\omega_{\rm{ef}}$ in the frame co-rotating with the drive frequency. For a transmon qutrit, the ratio of the Josephson energy $E_{J}$ to the charging energy $E_{C}$ is very large, and in the extreme limit $E_{J}/E_{C}\rightarrow\infty$, the transmon energy levels approach those of a harmonic oscillator for low-energy states, with nearly constant but slightly decreasing spacing due to residual anharmonicity. The qutrit is driven by a microwave field with frequency $\omega_{d}(t)$.
@@
We consider a rotating frame such that for the ground state $|\rm{g}\rangle$ the energy is shifted by $\Delta(t)$ and the zero of the energy is set at the first excited state $|\rm{e}\rangle$, where $\Delta(t)=\omega_{\rm ge}-\omega_{d}-(d\omega_{d}/dt)t$. The charging energy $E_{C}$ gives rise to the anharmonicity in the energy levels in superconducting qutrits such as the transmon and relates to the transition frequencies via $E_{\rm C}=\hbar\omega_{\rm{ge}}-\hbar\omega_{\rm{ef}}$. Thus, the detuning of the $|\mathrm{f}\rangle$ level is $\Delta(t)-E_{\rm C}/\hbar$. In this work, we set $E_{C}=0.3$ GHz, which is within the typical range for a transmon [Koch_2007]. The ratio of the Rabi frequencies of the two transitions is given by the ratio of the transition electric dipole moments. The dipole operator is proportional to the number operator $\hat{n}=-i\partial/\partial\varphi$ and thus the dipole matrix elements are $g_{n,n+m}\sim\langle n|\hat{n}|n+m\rangle$ and consequently, $g_{12}/g_{01}=\sqrt{2}$ and $g_{02}=0$.
@@
Combining these considerations, we arrive at the Hamiltonian in the rotating wave approximation, as detailed in Ref. [Bjorkman_2025]:
@@
Figure 1: The energy diagram of our frequency-modulation scheme shown for two different detuning functions, one linear (blue) and one quintic (green), for time $t\in[-T/2,T/2]$. The transition frequencies are $\omega_{\rm{ge}}$ and $\omega_{\rm{ef}}$; the modulation depth is denoted by $\Delta_{\rm max}$; and the frequency offset is $\delta$. The two-photon transition frequency $\omega_{\rm{gf}}^{\rm 2ph}$ is exactly halfway between $\omega_{\rm{ge}}$ and $\omega_{\rm{ef}}$.
2.2 Pulse models
@@
The shape of the drive pulse plays a central role in achieving efficient and robust state transfer. Several envelope functions have been explored in two-state and three-state quantum models, each with distinct advantages and limitations [Rosen1932, Demkov1969, Bambini1981, Allen1987, Carroll1986, Morris_1964, Rangelov_2005, Rangelov_2012, Yuan_2023]. Constant-amplitude pulses can drive transitions via Rabi oscillations but require precise timing and are highly sensitive to detuning errors [Rabi1937]. Landau-Zener transitions, which use a linearly time-dependent detuning with a constant coupling, offer analytical solutions but can suffer from unwanted population leakage [Landau1932, Zener1932]. Further, exponential pulses have been applied in time-dependent quantum transition models but often suffer from long tails, reducing efficiency by causing non-negligible residual population transfer outside the intended time window [Demkov1964, Nikitin1970].
@@
Smoothly varying pulses, such as Gaussian and hyperbolic secant envelopes, provide greater control over transition probabilities. Gaussian pulses minimize spectral side lobes and have been widely used in adiabatic passage schemes [Demirplak_2003, Vasilev2004, Vasilev_2005, Guerin_2011, Rangelov_2012, Yuan_2023]. However, achieving complete population transfer with Gaussian pulses requires precise detuning control. On the other hand, hyperbolic secant pulses are analytically solvable and naturally suited for robust adiabatic and nonadiabatic population transfer, making them widely used in RAP and related techniques [Rosen1932, Demkov1969, Bambini1981, Allen1987, Carroll1986, Ishkhanyan_2014, Ishkhanyan_2015].
@@
In this work, we focus on two pulse envelopes: generalized super-Gaussian and hyperbolic secant varieties. For appropriately chosen parameters, both envelope types are flat at the top, meaning the pulse amplitude remains approximately constant near the center of the drive, around $t=0$, where the avoided crossing typically occurs. This flat-top structure grants a greater degree of detuning robustness by maintaining a high Rabi frequency over the critical transition region, thereby extending the adiabatic window and improving population transfer even under moderate detuning errors [Kuzmanovic_2024]. The super-Gaussian envelope allows for greater control over the transition window by flattening the peak and steepening the decay, which can simultaneously be advantageous in terms of requiring lower amplitudes. Meanwhile, the hyperbolic secant envelope serves as a benchmark due to its well-established analytical properties. These envelope functions are defined as follows:
@@
for the super-Gaussian envelope and
@@
[…]
The envelope parameters that we seek to multi-objectively optimize are the amplitude $\Omega_{0}$, envelope order $n$, and the parameter $\beta$ (for super-Gaussian pulses) or $\epsilon$ (for hyperbolic secant pulses), which determine the envelope width and steepness. The envelope order $n$ generalizes the super-Gaussian envelopes in that if $n=2$ we recover a Gaussian profile, $n=4$ corresponds to the so-called super-Gaussian variety, and as $n\rightarrow\infty$ the pulse approaches a rectangular form. Amplitudes $\Omega_{0}$ are searched for in the range $[0,0.1]$ GHz, consistent with experimental constraints that help mitigate parasitic effects such as leakage into $|\rm{f}\rangle$ and higher states, as well as AC Stark shifts, which are known to degrade qubit coherence and fidelity in similar systems [Shillito_2022]. The envelope order $n$ is scanned over the range $(0,10]$, while $\beta$ and $\epsilon$ are searched within [6,0)[-6,0) and $(0,6]$, respectively.
Additionally, we consider four detuning models and assign each to the two envelope types. Our first detuning model is linear and the model name, when assigned to the super-Gaussian envelope, is SG1 and SECH1 when assigned to the hyperbolic secant envelope. For subsequent models, we hereafter adopt the prefixes SG and SECH for the model names.
@@
where the prefactor $k_{1}$ determines the slope of our detuning function and is a parameter we optimize by searching in the range $k_{1}\in[-0.3,0.3]$ GHz. Since Eq. (4) is a linear detuning function, the parameter $k_{1}$ is the modulation depth, which typically does not exceed half the charging energy, i.e., $E_{C}/2=0.15$ GHz in this work. Further, we center our modulation about the frequency offset $\delta$, i.e., by sweeping $(2\Delta_{\rm max}/T)t$ from t=T/2t=-T/2 to $t=T/2$.
We also explore a quintic detuning function,
$\Delta(t)=\sum\limits_{j=1}^{5}k_{j}(2t/T)^{j}+\delta,$ (5)
where $k_{j>1}$ are each searched for in the range [0.1,0.1][-0.1,0.1] GHz. For this detuning function, the effect of the parameter $k_{1}$ is the same as in Eq. (4); whereas, the prefactor for the quadratic term, $k_{2}$, makes the detuning function convex (upward-bending) when $k_{2}>0$ and concave (downward-bending) when $k_{2}<0$. Further, the parameter $k_{3}$, the cubic prefactor, introduces asymmetry and has the effect of making the right side ($t>0$) of the detuning function steeper and the left side ($t<0$) shallower when $k_{3}>0$ and vice versa when $k_{3}<0$. Still further, the quartic ($k_{4}$) and quintic ($k_{5}$) prefactors have similar effects to $k_{2}$ and $k_{3}$, respectively, but with stronger effects at larger $|\absolutevalue{t}$. Overall, the symmetry properties are governed by the parity of each term: the even-powered terms ($k_{2}$, $k_{4}$) contribute symmetric components about $t=0$, while the odd-powered terms ($k_{1}$, $k_{3}$, $k_{5}$) contribute antisymmetric components. A detuning function composed purely of odd-powered terms is antisymmetric about the origin, but adding even-powered terms breaks this antisymmetry, resulting in an overall asymmetric function. Thus, the inclusion of both even and odd terms enables finer control over both the shape and symmetry of the detuning profile.
We also examine detuning functions that incorporate the hyperbolic tangent. Motivated by the Allen–Eberly [Allen1987] and Hioe–Carroll [Carroll1986] models—both of which employ hyperbolic secant pulse envelopes—we respectively consider the following forms:
$\Delta(t)=k_{1}\tanh(2\gamma_{1}t/T)+\delta$ (6)